Two fair dice are tossed and the up face on each die is recorded. Find the probability of observing each of the following events: A:{the difference of the numbers is 1} B:{the sum of the numbers is equal to one} C:{the sum of the numbers is even} P(A)= P(B)= P(C)=
Answer #1

there are total 36 outcomes which is possible after tossed two fair dies and these are given as

(1,1),(1,2),(1,3),(1,4),(1,5),(1,6),

(2,1),(2,2),(2 ,3),(2,4),(2,5),(2,6),

(3,1),(3,2),(3,3),(3,4),(3,5 ),(3,6),

(4,1)(4,2)(4,3)(4,4)(4,5),(4,6),

(5,1),(5,2 ),(5,3)(5,4),(5,5),(5,6),

(6,1),(6,2),(6,3),(6,4),( 6,5),(6,6)

A) possible no of outcomes whose pair of difference is 1 is 5

P{the difference of the numbers is 1} = 5/36

B) possible numbers of outcomes in pairs whose sum of the numbers is equal to one is 0

so P{the sum of the numbers is equal to one}= 0

C)possible numbers of outcomes in pair whose  sum of the numbers is even is 16

P(the sum of the numbers is even) = 16/36 =4/9